OpenStudy (anonymous):

Gino designed a garden for his rectangular backyard. The design he drew on paper is ten inches long. The length of his backyard is forty feet. If his drawing is to scale, and his backyard is twenty-eight feet wide, how many inches wide is his drawing? Write a proportion using words. Then, use the proportion to calculate the drawing width. Show your work

OpenStudy (anonymous):

@jigglypuff314

jigglypuff314 (jigglypuff314):

same idea as last time :) the ratio of the lengths is equal to the ratio of the widths so 10 / 40 = ? / 28 solve for the ?

OpenStudy (anonymous):

i dont really get how to do this stuff

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

what is number is the question mark supposed to be

jigglypuff314 (jigglypuff314):

10 / 40 = x / 28 40x = 280 <- can you solve for x from there?

OpenStudy (anonymous):

yes I still have like 2 questions